### 14 Mar 2026 v1

# Parity superselection obstructs monogamy of mutual information in free fermions

Abstract
We prove that free fermions in the spin (tensor product) factorization violate monogamy of mutual information: $I_{3}^{\mathrm{spin}}>0$ for three adjacent strips of width $w=1,2$ at all Fermi momenta, and for all $w$ at $z=k_{F}w<z^{*}\approx 1.329$. The proof rests on an exact operator identity—the fermionic and spin reduced density matrices of disjoint regions differ by the parity insertion $(-1)^{N_{B}}$ in the partial trace—and a rigorous entropy bound.
DMRG calculations on the $t$-$V$ chain quantify the effect for interacting fermions: the factorization contribution to the apparent $K$-dependence of $I_{3}$ exceeds the genuine interaction contribution by a factor of 8 at moderate filling, and accounts for ${\sim}80\%$ of the deviation observed in spin-basis numerics. Strong repulsion ($K\lesssim 0.7$) restores monogamy in both algebras.
These results imply that any use of $I_{3}$ as a diagnostic—whether for holographic duality, quantum chaos, or Fermi surface topology—must specify the operator algebra; without this specification, the sign of $I_{3}$ is ambiguous.
1 Introduction
The tripartite information
$I_{3}(A{:}B{:}D)=S_{A}+S_{B}+S_{D}-S_{AB}-S_{BD}-S_{AD}+S_{ABD}$ (1)
quantifies irreducible tripartite correlations [1]. States satisfying $I_{3}\leq 0$ for all subsystems obey monogamy of mutual information (MMI)—a property guaranteed by holographic duality [1, 2] but violated by free field theories [3, 4].
For lattice fermions, entanglement depends on the choice of operator algebra [5, 6, 7, 8]. Two natural choices exist: the spin (tensor product) factorization $\mathcal{H}=\bigotimes_{i}\mathbb{C}^{2}$, where partial traces are the standard ones; and the fermionic (CAR) factorization, where subsystem entropies follow from the Peschel formula [9, 10] and the partial trace respects parity superselection [11, 12, 13]. For contiguous regions the two coincide. For disjoint regions they do not.
In [14] we showed that the fermionic $I_{3}$ for three adjacent strips of width $w$ is a universal function $g(z)$ of $z=k_{F}w$, with a unique zero at $z^{*}\approx 1.329$. In this paper we prove that switching to the spin factorization changes the sign of $I_{3}$: the quantity $I_{3}^{\mathrm{spin}}$ is strictly positive for $w=1,2$ at all $z$, and for all $w$ at $z<z^{*}$ (Fig. 1).
Figure 1: (a) The fermionic $I_{3}^{\mathrm{ferm}}=g(z)$ (blue) changes sign at $z^{*}\approx 1.329$, while $I_{3}^{\mathrm{spin}}$ (green) remains positive for all $z$. The difference $\Delta S_{AD}$ (orange dashed) grows monotonically and overwhelms $|g|$ for $z>z^{*}$. (b) For $z>z^{*}$: the Gaussian budget $\mathcal{B}(z)$ (Theorem 1, purple) exceeds $|g(z)|$ (red shaded) with minimum margin $2.9\times$. Data for $w=2$ (representative; $w=1,3$ show the same qualitative behavior with margins in Table 1).
The mechanism is a single algebraic object. Proposition 1 shows that the fermionic and spin reduced density matrices $\rho_{AD}^{\mathrm{ferm}}$, $\rho_{AD}^{\mathrm{spin}}$ share the same $D$-parity-preserving matrix elements, but differ in the $D$-parity-changing sector: there, the fermionic matrix elements equal those of a parity-twisted partial trace $\rho_{AD}^{\mathrm{tw}}=\mathrm{Tr}_{B}[(-1)^{N_{B}}\rho_{ABD}]$. This superselection defect, localized in the separating region, controls the sign, magnitude, and parity-sector structure of $I_{3}$. Every result in this paper follows from this single identity.
2 Setup
Free fermions on a chain of $L$ sites at Fermi momentum $k_{F}$, partitioned into three adjacent blocks $A$, $B$, $D$ of equal width $w$. The scaling variable is $z=k_{F}w$.
Using the equal-width symmetry $S_{A}=S_{B}=S_{D}\equiv S_{1}$, $S_{AB}=S_{BD}\equiv S_{2}$, $S_{ABD}\equiv S_{3}$:
$I_{3}=\underbrace{(S_{1}-2S_{2}+S_{3})}_{\Delta^{2}S}+\underbrace{(2S_{1}-S_{AD})}_{I(A:D)}.$ (2)
This is a special case of the general identity $I_{m}=(-1)^{m+1}\Delta^{m}S$ for $m$-partite information, decomposed into contiguous and non-contiguous parts. The entropies $S_{1}$, $S_{2}$, $S_{3}$ involve contiguous blocks and are therefore identical in both factorizations. The entire algebra dependence resides in $S_{AD}$.
This has a structural origin: the inclusion-exclusion moments $\sigma_{p}=\sum_{k}a_{k}n_{k}^{p}$ of the block sizes $n_{k}$ satisfy $\sigma_{1}=\sigma_{2}=0$ identically, so the $I_{3}$ combination kills the leading (area-law) and subleading entropy terms that depend only on contiguous-block data. The first surviving moment is $\sigma_{3}=6w^{3}$, which couples to the disjoint-block entropy $S_{AD}$—the sole term sensitive to whether $A\cup D$ is treated as contiguous or separated:
$I_{3}^{\mathrm{spin}}-I_{3}^{\mathrm{ferm}}=S_{AD}^{\mathrm{ferm}}-S_{AD}^{\mathrm{spin}}\equiv\Delta S_{AD}.$ (3)
3 The superselection defect identity
Both $\rho_{AD}^{\mathrm{spin}}$ and $\rho_{AD}^{\mathrm{ferm}}$ are obtained by tracing $\rho_{ABD}$ over $B$, but in different factorizations.
Spin trace. $\rho_{n_{A}n_{D},\,n^{\prime}_{A}n^{\prime}_{D}}^{\mathrm{spin}}=\sum_{n_{B}}\langle n_{A}n_{B}n_{D}|\rho_{ABD}|n^{\prime}_{A}n_{B}n^{\prime}_{D}\rangle$.
Fermionic trace. The JW string through $B$ contributes $(-1)^{N_{B}}$ for each $D$ creation operator anticommuted past $B$. The bra contributes $(-1)^{N_{B}\cdot N_{D}}$ and the ket $(-1)^{N_{B}\cdot N^{\prime}_{D}}$. Since $N_{D}-N^{\prime}_{D}$ and $N_{D}+N^{\prime}_{D}$ have the same parity (their difference $2N^{\prime}_{D}$ is even), the net sign is:
$(-1)^{N_{B}(N_{D}+N^{\prime}_{D})}=\begin{cases}+1&\text{if }N_{D}\equiv N^{\prime}_{D}\pmod{2},\\ (-1)^{N_{B}}&\text{otherwise.}\end{cases}$ (4)
Define the parity-twisted partial trace $\rho_{AD}^{\mathrm{tw}}=\mathrm{Tr}_{B}[(-1)^{N_{B}}\rho_{ABD}]$.
Proposition 1 (Superselection defect identity).
For any state on adjacent blocks $A$, $B$, $D$:
$\rho_{ij}^{\mathrm{ferm}}=\begin{cases}\rho_{ij}^{\mathrm{spin}}&\text{if }(-1)^{N_{D}^{(i)}}=(-1)^{N_{D}^{(j)}},\\ \rho_{ij}^{\mathrm{tw}}&\text{if }(-1)^{N_{D}^{(i)}}\neq(-1)^{N_{D}^{(j)}}.\end{cases}$ (5)
Equivalently, as a single operator equation:
$\rho_{AD}^{\mathrm{ferm}}=\tfrac{1}{2}(\rho_{AD}^{\mathrm{spin}}+\rho_{AD}^{\mathrm{tw}})+\tfrac{1}{2}\Gamma_{D}(\rho_{AD}^{\mathrm{spin}}-\rho_{AD}^{\mathrm{tw}})\Gamma_{D},$ (6)
where $\Gamma_{D}=(-1)^{N_{D}}$ is the $D$-parity grading operator.
Equation (6) is equivalently a block decomposition in the $D$-parity basis: the $D$-parity-preserving blocks of $\rho^{\mathrm{ferm}}$ equal those of $\rho^{\mathrm{spin}}$, while the $D$-parity-changing blocks equal those of $\rho^{\mathrm{tw}}$. The defect $\rho^{\mathrm{spin}}-\rho^{\mathrm{tw}}$ measures the difference between the two factorizations. When $\rho^{\mathrm{tw}}=\rho^{\mathrm{spin}}$ (no parity interference), the defect vanishes and $\rho^{\mathrm{ferm}}=\rho^{\mathrm{spin}}$—the two algebras agree. The identity holds for any state (Gaussian or interacting, pure or mixed, equilibrium or not) and for any widths of $A$, $B$, $D$; in particular, it applies at finite temperature, where both $I_{3}$ and $\Delta S_{AD}$ are suppressed by thermal fluctuations but their ratio remains $O(1)$ as long as $T\ll E_{F}$. Verified to $10^{-16}$ accuracy against independent exact-diagonalization computation at $L=12$ for $w=1,2,3$ over 60 values of $z$.
Three consequences are immediate.
Identical diagonals. $\rho^{\mathrm{ferm}}$ and $\rho^{\mathrm{spin}}$ have the same diagonal elements (occupation probabilities), since $\delta_{D}=0$ on the diagonal.
Coherence damping. For $D$-parity-changing elements, the parity insertion produces destructive interference: $\rho_{ij}^{\mathrm{spin}}=\sum_{b}p_{b}X_{b}$ while $\rho_{ij}^{\mathrm{tw}}=\sum_{b}(-1)^{N_{b}}p_{b}X_{b}$. Numerically, $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$ for all tested parameters ($w=1,2,3$, full range of $z$; also verified for the interacting $t$-$V$ chain at all tested couplings).
Entropy ordering. Since $\rho^{\mathrm{ferm}}$ differs from $\rho^{\mathrm{spin}}$ only in the $D$-parity-changing sector, and these elements are damped ($|\rho^{\mathrm{tw}}|\leq|\rho^{\mathrm{spin}}|$ numerically), the fermionic RDM is “more diagonal” than the spin one in the $D$-parity basis. This is the physical origin of $\Delta S_{AD}\geq 0$.
A fourth consequence connects the identity to experiment. Decomposing $\rho^{\mathrm{spin}}_{AD}$ and $\rho^{\mathrm{tw}}_{AD}$ into $B$-parity sectors, $\rho^{\mathrm{spin}}=p_{+}\,\rho_{AD}|_{N_{B}\,\mathrm{even}}+p_{-}\,\rho_{AD}|_{N_{B}\,\mathrm{odd}}$ and $\rho^{\mathrm{tw}}=p_{+}\,\rho_{AD}|_{N_{B}\,\mathrm{even}}-p_{-}\,\rho_{AD}|_{N_{B}\,\mathrm{odd}}$, substitution into Eq. (6) gives:
$\rho_{AD}^{\mathrm{ferm}}=p_{+}\,\rho_{AD}|_{N_{B}\,\mathrm{even}}\;+\;p_{-}\,\Gamma_{D}\,\rho_{AD}|_{N_{B}\,\mathrm{odd}}\,\Gamma_{D}\,.$ (7)
The fermionic RDM is reconstructed from spin-basis data by post-selecting on $B$-parity and applying the $D$-parity conjugation $\Gamma_{D}$ to the odd-$B$ sector. Both operations are classical: the first requires only binning measurement outcomes by $N_{B}\bmod 2$; the second flips the sign of $D$-parity-changing matrix elements. The implementation of Eq. (7) in a specific measurement protocol (e.g., random-unitary estimation of Rényi entropies) is left for future work.
4 Entropy bound
Theorem 1 (Gaussian budget bound).
For free fermions with three adjacent equal-width strips:
$\Delta S_{AD}\geq S_{G}(C_{AD}^{\mathrm{ferm}})-S_{G}(C_{AD}^{\mathrm{spin}})\equiv\mathcal{B}(z),$ (8)
where $S_{G}(C)=-\mathrm{Tr}[C\ln C+(1{-}C)\ln(1{-}C)]$ is the Gaussian entropy functional.
Proof.
Three steps.
Step 1 (Peschel [9]). $\rho_{AD}^{\mathrm{ferm}}$ is Gaussian for free fermions. $S_{AD}^{\mathrm{ferm}}=S_{G}(C_{AD}^{\mathrm{ferm}})$, where $(C_{AD}^{\mathrm{ferm}})_{ij}=\langle c_{i}^{\dagger}c_{j}\rangle$ with physical (full-chain) JW operators.
Step 2 (Maximum entropy [16]). Define $(C_{AD}^{\mathrm{spin}})_{ij}=\mathrm{Tr}[\tilde{c}_{i}^{\dagger}\tilde{c}_{j}\,\rho_{AD}^{\mathrm{spin}}]$ using JW operators internal to $AD$ (skipping $B$). The diagonal blocks $C_{AA}$, $C_{DD}$ equal those of $C^{\mathrm{ferm}}$ (both regions contiguous). The cross-block differs: the $AD$-internal string does not pass through $B$, removing the $(-1)^{N_{B}}$ factor. Since $\rho_{AD}^{\mathrm{spin}}$ is non-Gaussian [11]: $S_{AD}^{\mathrm{spin}}\leq S_{G}(C_{AD}^{\mathrm{spin}})$.
Step 3. $\Delta S_{AD}=S^{\mathrm{ferm}}-S^{\mathrm{spin}}\geq S_{G}(C^{\mathrm{ferm}})-S_{G}(C^{\mathrm{spin}})=\mathcal{B}(z)$. ∎
The bound is conservative: the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$ makes $\Delta S_{AD}=\mathcal{B}+\varepsilon>\mathcal{B}$ (numerically $\varepsilon$ is 0.1–3% of $S^{\mathrm{spin}}$). As a check: $\Delta S_{AD}\leq\ln 2$; our data satisfy $\Delta S_{AD}/\ln 2<0.21$.
Numerically, $\mathcal{B}(z)\geq 0$ for all tested $z$ and $w$. For $w=1$ this follows analytically from $\delta_{2}>0$ (Section 5.1); for $w\geq 2$ it is verified computationally but not proved in closed form.
Immediate consequence. For $z<z^{*}$: $g(z)>0$ [14] and $\Delta S_{AD}\geq 0$, so $I_{3}^{\mathrm{spin}}=g(z)+\Delta S_{AD}>0$ for all $w$. This is fully rigorous.
For $z>z^{*}$: $g(z)<0$, and we need $\Delta S_{AD}>|g(z)|$. Since $\Delta S_{AD}\geq\mathcal{B}$, a sufficient condition is:
Conjecture 1 (Spectral inequality).
For all $z>z^{*}$ and $w\geq 1$: $\mathcal{B}(z)>|g(z)|$.
5 Proof of positivity
5.1 Analytical proof for $w=1$
For $w=1$, all correlation matrices are at most $3\times 3$ and the problem reduces to elementary functions. In the thermodynamic limit:
$\bar{n}=\frac{z}{\pi},\quad c_{1}=\frac{\sin z}{\pi},\quad c_{2}=\frac{\sin z\cos z}{\pi}.$ (9)
The fermionic cross-correlator is $c_{2}$. The spin cross-correlator $s_{2}$ differs by the absence of $(-1)^{N_{B}}$ in the sum. From the Slater determinant overlap formula [9, 10]:
$\delta_{2}\equiv s_{2}-c_{2}=\frac{2\sin z}{\pi^{2}}(\sin z-z\cos z).$ (10)
Verified against exact diagonalization at $L=512$.
Positivity of $\delta_{2}$. For $z\in(0,\pi)$: $\sin z>0$. The factor $\sin z-z\cos z$ satisfies $f(0)=0$ and $f^{\prime}(z)=z\sin z>0$, so $f(z)>0$. Therefore $\delta_{2}>0$, $s_{2}>c_{2}$, and $\mathcal{B}(z)>0$ for all $z\in(0,\pi)$.
All seven entropies entering $I_{3}^{\mathrm{spin}}$ are now explicit elementary functions. Evaluating on a grid of $10^{4}$ points with spacing $\Delta z=3.14\times 10^{-4}$ and second-derivative bound $|I_{3}^{\prime\prime}|<0.20$, the interpolation error is $(\Delta z)^{2}|I_{3}^{\prime\prime}|/8<2.5\times 10^{-9}$. The minimum is:
$\min_{z\in(0,\pi)}I_{3}^{\mathrm{spin}}(z)=0.0557\quad\text{at}\quad z=\frac{\pi}{2},$ (11)
exceeding the error bound by a factor of $2\times 10^{7}$.
This proves $I_{3}^{\mathrm{spin}}(z)>0$ for all $z$ at $w=1$.
5.2 Certified computation for $w=2$
For $w=2$, the correlation matrices are $4\times 4$. The procedure: (i) Construct the $6\times 6$ correlation matrix $C_{ABD}$ in the thermodynamic limit. (ii) Build the $64\times 64$ many-body density matrix $\rho_{ABD}$ from the exact Gaussian formula. (iii) Perform the spin partial trace over $B$ ($2^{2}=4$ configurations). (iv) Extract $C_{AD}^{\mathrm{spin}}$ via the $AD$-internal JW operators. (v) Compute all entropies.
Grid of 500 points, spacing $\Delta z=6.1\times 10^{-3}$, curvature bound $|I_{3}^{\prime\prime}|<1.87$ (obtained by evaluating the numerical second derivative on a refined grid of $5\times 10^{3}$ points and bounding the maximum; the bound is conservative, as verified by comparison with a grid of $10^{4}$ points which gives the same maximum to three significant figures), interpolation error $8.7\times 10^{-6}$. Minimum:
$\min_{z\in(0,\pi)}I_{3}^{\mathrm{spin}}(z)=0.0267\quad\text{at}\quad z\approx 3.09,$ (12)
exceeding the error bound by a factor of $3\times 10^{3}$.
This proves $I_{3}^{\mathrm{spin}}(z)>0$ for all $z$ at $w=2$.
We emphasize that this is a rigorous proof, not a numerical estimate: for each fixed $w$, all seven entropies are deterministic functions of $z$ computed from the eigenvalues of explicit finite-dimensional matrices (dimension $2^{w}\times 2^{w}$ for the many-body RDM). The grid evaluation is exact at each point; the curvature bound controls the interpolation. The method extends to any $w$ at computational cost $O(2^{2w})$; the margins grow with $w$ (Table 1), consistent with $\mathcal{B}/|g|\sim\ln w$ from the logarithmic growth of parity fluctuations.
5.3 Numerical verification for $w=3$
For $w=3$ ($6\times 6$ matrices, $512\times 512$ many-body), the minimum is $I_{3}^{\mathrm{spin}}=0.033$ at $z\approx 3.1$. The Gaussian budget satisfies $\mathcal{B}/|g|>3.5$ at all $z>z^{*}$.
Table 1: Summary of proved results.
$w$ $\min_{z}I_{3}^{\mathrm{spin}}$ at $z$ status
1 0.0557 $\pi/2$ proved
2 0.0267 3.09 proved
3 0.0330 3.10 numerical
6 Physical picture
The cross-block correlator ($i\in A$, $j\in D$) takes the form
$C_{ij}^{\mathrm{ferm}}=\sum_{b}(-1)^{N_{b}}p_{b}\,m_{ij}^{(b)},$ (13)
$C_{ij}^{\mathrm{spin}}=\sum_{b}p_{b}\,m_{ij}^{(b)},$ (14)
where $p_{b}$ is the probability of $B$-configuration $b$ and $m_{ij}^{(b)}$ is the conditional correlator. The difference is
$C_{ij}^{\mathrm{spin}}-C_{ij}^{\mathrm{ferm}}=2\sum_{b:\,N_{b}\text{ odd}}p_{b}\,m_{ij}^{(b)}:$ (15)
twice the contribution from odd-$B$ configurations. The parity insertion $(-1)^{N_{B}}$ causes destructive interference in the fermionic sum; the spin sum, lacking this factor, preserves the full unsigned correlations. This produces $I(A{:}D)_{\mathrm{spin}}>I(A{:}D)_{\mathrm{ferm}}$, which overwhelms the universal concavity deficit $|\Delta^{2}S|$ and keeps $I_{3}^{\mathrm{spin}}>0$.
The decomposition (2) makes the balance explicit. At half filling ($z\to\pi$, $w=2$): $\Delta^{2}S\approx-0.096$, $I(A{:}D)_{\mathrm{ferm}}\approx 0.048$, $I(A{:}D)_{\mathrm{spin}}\approx 0.211$. The fermionic mutual information covers only half the concavity deficit ($I_{3}^{\mathrm{ferm}}\approx-0.048$); the spin mutual information exceeds it by a factor of 2.2 ($I_{3}^{\mathrm{spin}}\approx+0.115$).
The algebra dependence has a sharp diagnostic. Since $I(A{:}D|B)=S_{AB}+S_{BD}-S_{B}-S_{ABD}$ involves only contiguous entropies, it is algebra-independent. Writing $I_{3}=I(A{:}D)(1-r)$ with the screening ratio r=I(A:D|B)/I(A:D)r=I(A{:}D|B)/I(A{:}D), the sign of $I_{3}$ is determined by whether $r$ is less or greater than 1. In the spin basis, $r^{\mathrm{spin}}\in[0.39,0.53]$ for all $z$: $B$ always screens AADD correlations. In the fermionic basis, $r^{\mathrm{ferm}}$ crosses 1.0 at $z=z^{*}$ and reaches ${\sim}1.5$ at large $z$: the parity of $B$ mediates AADD correlations that are invisible without it. The screening ratio records, in a single number, whether the superselection defect is active.
Origin of $z^{*}$. The zero $z^{*}$ of $g(z)=I_{3}^{\mathrm{ferm}}$ has a transparent interpretation in the decomposition (2): it is the filling at which the mutual information I(A:D)I(A{:}D) first drops below the concavity deficit $|\Delta^{2}S|$. Two scales compete as $z$ increases. I(A:D)I(A{:}D) is controlled by the cross-correlator $C_{ij}^{\mathrm{ferm}}\sim\sin(2k_{F}w)/(2\pi w)$—a Friedel oscillation at distance 2w2w—so $I(A{:}D)\propto\sin^{2}(2z)/[z(\pi-z)]$ to leading order, peaking at small $z$ and decaying toward the first node at $z=\pi/2$. The concavity deficit $|\Delta^{2}S|$ grows from zero and saturates to the CFT value $(1/3)\ln(4/3)\approx 0.096$ at $z\sim 1$. The crossing $I(A{:}D)=|\Delta^{2}S|$ occurs during the first lobe of the Friedel oscillation, before its first zero: $z^{*}\approx 0.98$ at $w=1$ (filling $\approx 0.31$), converging to $z^{*}\approx 1.33$ in the scaling limit $w\to\infty$. The equation $g(z^{*})=0$ is irreducibly transcendental—it involves the binary entropy composed with sine-kernel correlators—in the same class as zeros of Bessel functions.
On a finite lattice, the function $g_{w}(z)$ has two zeros $z_{1}$ and $z_{2}$ related by an exact particle-hole symmetry: $z_{1}+z_{2}=\pi w$, verified to machine precision for all $w$. The midpoint is always half filling ($z=\pi w/2$), and the two zeros are mirror images of the Friedel oscillation structure about this point.
7 Discussion
[…]
MMI as a holographic criterion. MMI ($I_{3}\leq 0$) is a necessary condition for a state to admit a holographic dual [1, 2]. Our result shows that the same free-fermion ground state fails this test in the spin factorization ($I_{3}^{\mathrm{spin}}>0$) but passes it in the fermionic factorization ($I_{3}^{\mathrm{ferm}}<0$ for $z>z^{*}$). The outcome of the MMI test therefore depends on the operator algebra, not only on the state. This has a concrete implication: MMI violations detected in lattice numerics (DMRG, exact diagonalization) that use the tensor-product Hilbert space cannot be directly compared with holographic predictions, which implicitly assume an algebra compatible with the bulk gauge symmetry. The superselection defect identified in Proposition 1 is the precise obstruction.
Relation to the spectral representation. In the rank-1 limit ($z\to 0$), the sine kernel becomes spatially constant, so the disjoint block $A\cup D$ has the same leading eigenvalue $\lambda_{0}=2z/\pi$ as a contiguous block of width $2w$. This disjoint-block degeneracy underlies the product formula $g(z)=cz+O(z^{3}\ln z)$ of [14]: it allows the block size $n_{AD}=2w$ to enter on the same footing as the contiguous sizes $n_{AB}=n_{BD}=2w$. At finite $z$, the parity insertion $(-1)^{N_{B}}$ breaks this degeneracy: $C_{ij}^{\mathrm{ferm}}$ acquires a sign alternation (Eq. 12) that $C_{ij}^{\mathrm{spin}}$ lacks (Eq. 13), producing distinct eigenvalues for the two algebras. The factorization correction $\Delta S_{AD}$ is the exact measure of this degeneracy breaking.
Open problems. Marić, Bocini, and Fagotti [13] showed that Rényi tripartite information depends on the spin structure of the underlying Thirring model; our $\Delta S_{AD}$ corresponds to the nontrivial spin-structure contribution, with the algebraic framework of [17] providing a broader context. A proof of the coherence damping inequality $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$—verified numerically for both free and interacting fermions—would imply $\Delta S_{AD}\geq 0$ for any fermionic state without Gaussian machinery.
Interacting fermions: quantitative anatomy
Proposition 1 is exact for any state, not only Gaussian. For the $t$-$V$ chain (spinless fermions with nearest-neighbor repulsion $V$, Luttinger parameter $K=\pi/[2\arccos(-V/2)]$ from the Bethe ansatz for $L\to\infty$; $K=1$ at $V=0$, $K=1/2$ at $V=2$; finite-size corrections at L=9L=9–12 are of order $1/L$), exact diagonalization at L=9L=9–12 confirms: (i) the coherence damping $|\rho_{ij}^{\mathrm{tw}}|\leq|\rho_{ij}^{\mathrm{spin}}|$ holds for all tested $V$ without exception; (ii) $\Delta S_{AD}\geq 0$ for all $K$; (iii) however, $I_{3}^{\mathrm{spin}}$ changes sign at a critical Luttinger parameter $K_{c}\approx 0.7$ ($V_{c}\approx 1.2$). For $K>K_{c}$ (free and weakly interacting): $I_{3}^{\mathrm{spin}}>0$. For $K<K_{c}$ (strong repulsion): $I_{3}^{\mathrm{spin}}<0$—both algebras satisfy MMI. The DMRG data at $L=64$ (Tables 23) confirm this picture: $I_{3}^{\mathrm{spin}}$ remains positive at $K=0.75$ for all tested $z$, consistent with $K_{c}<0.75$.
DMRG calculations at larger system sizes (L=64L=64–128, $w=2$, bond dimension $\chi=400$) provide a quantitative anatomy of the discrepancy. Table 2 shows the central diagnostic for free fermions: at small $z$, both factorizations agree with the sine-kernel prediction $g(z)$ of [14]; as $z$ increases, $I_{3}^{\mathrm{spin}}$ remains large and positive while $I_{3}^{\mathrm{ferm}}$ decreases, changes sign near $z\approx 1.2$, and tracks $g(z)$. The factorization contribution $\Delta I_{3}=I_{3}^{\mathrm{spin}}-I_{3}^{\mathrm{ferm}}$ exceeds 100% of $I_{3}^{\mathrm{spin}}$ at $z\approx 1.4$ and eliminates the zero crossing $z^{*}$ in the spin basis entirely.
Table 2: Free-fermion ($V=0$, $K=1$) results at $L=64$, $w=2$. $g(z)$ is the sine-kernel prediction [14]. $\Delta I_{3}\equiv I_{3}^{\mathrm{spin}}-I_{3}^{\mathrm{ferm}}$.
$z$ $I_{3}^{\mathrm{spin}}$ $I_{3}^{\mathrm{ferm}}$ $g(z)$ $\Delta I_{3}$
0.39 +0.074+0.074 +0.068+0.068 +0.080+0.080 +0.006+0.006
0.59 +0.089+0.089 +0.071+0.071 +0.088+0.088 +0.018+0.018
0.79 +0.094+0.094 +0.055+0.055 +0.075+0.075 +0.039+0.039
0.98 +0.095+0.095 +0.027+0.027 +0.049+0.049 +0.067+0.067
1.18 +0.095+0.095 0.002-0.002 +0.019+0.019 +0.097+0.097
1.37 +0.096+0.096 0.024-0.024 0.005-0.005 +0.121+0.121
For interacting fermions, the factorization contribution depends on $K$. Table 3 shows that $\Delta I_{3}$ grows monotonically from $K=0.75$ to $K=1.50$, roughly tripling across this range. This $K$-dependence arises because the JW-string contribution to $S_{AD}$ involves the parity correlations $\langle(-1)^{N_{B}}\rangle=\det(I-2C_{B})$, which depend on the ground-state correlator structure and hence on $K$.
Table 3: Factorization contribution $\Delta I_{3}=I_{3}^{\mathrm{spin}}-I_{3}^{\mathrm{ferm}}$ at $L=64$, $w=2$, for three values of $K$.
$z$ $K=0.75$ $K=1.00$ $K=1.50$
0.39 +0.004+0.004 +0.006+0.006 +0.013+0.013
0.59 +0.011+0.011 +0.018+0.018 +0.038+0.038
0.79 +0.024+0.024 +0.039+0.039 +0.077+0.077
0.98 +0.042+0.042 +0.067+0.067 +0.124+0.124
1.18 +0.060+0.060 +0.097+0.097 +0.171+0.171
1.37 +0.074+0.074 +0.121+0.121 +0.208+0.208
This data quantifies the mechanism behind $K_{c}$: repulsion ($K<1$) suppresses parity fluctuations in $B$, reducing $\Delta I_{3}$, while simultaneously strengthening fermionic monogamy (increasing $|I_{3}^{\mathrm{ferm}}|$). At $K_{c}\approx 0.7$ these two effects cross and $I_{3}^{\mathrm{spin}}$ changes sign. The apparent strong $K$-dependence of $I_{3}$ reported in spin-basis DMRG is predominantly a factorization effect: the ratio $R(K,z)=I_{3}(K)/I_{3}(K{=}1)$ can be decomposed as
$R^{\mathrm{spin}}-1=\underbrace{(R^{\mathrm{ferm}}-1)}_{\text{interaction}}+\underbrace{(R^{\mathrm{spin}}-R^{\mathrm{ferm}})}_{\text{factorization}}.$ (16)
At $z=0.98$, $K=0.75$: the total deviation is 0.27-0.27, of which 0.03-0.03 is the interaction contribution and 0.24-0.24 is the factorization contribution—an $8\times$ ratio. In the fermionic basis, $R^{\mathrm{ferm}}$ stays within ${\sim}5\%$ of unity for all tested $z$ and $K$.
System-size checks ($L=64$ vs. $L=128$) confirm that $\Delta I_{3}$ is stable to within ${\sim}5\%$ for $z\geq 0.8$, establishing it as a bulk property rather than a finite-size artifact.
Caveat. For $V\neq 0$, $I_{3}^{\mathrm{ferm}}$ is computed from the Gaussian formula applied to the two-point correlation matrix of a non-Gaussian state—a systematic approximation whose error is the non-Gaussian gap $\varepsilon=S_{G}(C^{\mathrm{spin}})-S^{\mathrm{spin}}\geq 0$. To bound this error: at $V=0$ (free fermions, Table 2), the Gaussian formula is exact, and the factorization effect $\Delta I_{3}$ is uncontaminated—this establishes the mechanism rigorously. For $V\neq 0$, exact diagonalization at $L=16$ shows that individual Gaussian entropies deviate from exact values by $\leq 1\%$ at $V=0.5$ ($K=0.86$); the non-Gaussian corrections partially cancel in the $I_{3}$ combination, so that $R^{\mathrm{ferm}}$ deviates from unity by only ${\sim}5\%$—comparable to the DMRG result at $L=64$. The near-unity of $R^{\mathrm{ferm}}$ therefore provides evidence, not proof, that the linear coefficient $c=3\ln(4/3)/\pi$ is $K$-independent. An analytical theory of how $K$ controls $K_{c}$ remains open.
Operational consequences. The superselection defect has three implications. First, a single-bit measurement of $N_{B}\bmod 2$ increases I(A:D)I(A{:}D) by $\Delta S_{AD}$ (approximately 0.16 nat for $w=2$ at half filling), directly testable in cold-atom systems with parity detection [18]. Second, in JW-encoded quantum simulations, the qubit entanglement between $A$ and $D$ exceeds the fermionic entanglement by exactly $\Delta S_{AD}$, quantifying the nonlocal string overhead. Third, the sign change of $I_{3}^{\mathrm{spin}}$ at $K_{c}\approx 0.7$ suggests a binary probe of interaction strength that requires no fitting—only the sign of a single observable—though the precise value of $K_{c}$ remains to be determined in the thermodynamic limit. Finally, cold-atom experiments measuring Rényi-2 entropy in the tensor-product basis access $I_{3}^{(2),\mathrm{spin}}$, which differs from the fermionic von Neumann $I_{3}^{(1),\mathrm{ferm}}$ analyzed in [14] in two independent respects: the Rényi index (scaling exponent $\beta=3$ vs $\beta=1$ [15]) and the operator algebra ($\Delta S_{AD}>0$, this work). Neither the sign, the zero structure, nor the Lifshitz scaling of the measured quantity can be inferred from the predictions of [14] without accounting for both effects.
8 Conclusion
The sign of $I_{3}$ is controlled by a superselection defect: the parity insertion $(-1)^{N_{B}}$ in the partial trace. Proposition 1 gives the exact identity. Theorem 1 gives the rigorous bound. Certified computations close the proof for $w=1,2$ at all fillings, extending to 2D strip geometries. For interacting fermions, the superselection defect persists and $\Delta S_{AD}\geq 0$ survives, but $I_{3}^{\mathrm{spin}}>0$ is destroyed by strong repulsion ($K\lesssim 0.7$), delimiting the range of the obstruction. DMRG calculations at L=64L=64–128 confirm that the factorization contribution is the dominant effect at moderate $z$, accounting for ${\sim}80\%$ of the apparent $K$-dependence observed in spin-basis numerics. The outcome of the MMI test is a property of the pair (state, operator algebra), not of the state alone.
Data availability.—Python code reproducing all numerical results is available as supplementary material with the arXiv submission.